Star A on apparent magnitude of 7, and its light intensity is 100 times dimmer than that of star B, what is the apparent visual magnitude of star B?


1
Expert's answer
2021-01-26T08:41:12-0500

From the problem statement we know

"m_A=7"

"I_B=100I_A"

Then from Pogson's formula

"\frac{I_B}{I_A}=2.512^{m_A-m_B}"

write down


"\frac{100I_A}{I_A}=2.512^{7-m_B}"

"100=2.512^{7-m_B}"

"\lg{(100)}=\lg{(2.512^{7-m_B})}"

"2=0.4(7-m_B)"

"5=7-m_B"

"m_B=7-5"

"m_B=2"